40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1766     Entry time: Tue Nov 8 18:22:07 2016
Author: yinzi 
Type: DailyProgress 
Category: TempCtrl 
Subject: driver circuit update 

So after testing out the LT1012 op amp in many simple configurations and not having any of them work (and the measurements were the same whether they were powered or not), I decided to dig around for another op amp chip, which did work, so I think all 4 of LT1012 chips that I got from the 40m stock are just faulty (not sure what that means for the rest of the chips there).

After swapping out chips, the circuit worked functionally. Just from playing around with the function generator and oscilloscope, it looks like the gain is 0.5 around 20Hz (should be dropping off faster), but I'll try to get an actual frequency response measurement next time.

Also, the legs on the capacitors are really short and I'm not sure they're all actually making contact, so I'll probably extend them with some wire to be sure, but that might also be part of it.

ELOG V3.1.3-